OpenStudy (anonymous):

simplify 12+65divided by 5+(8divided by 4)

OpenStudy (anonymous):

27

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

=12+(65/5)+(8/4) =12+13+2 =27

OpenStudy (anonymous):

ope did it correctly